Thursday, April 16, 2009

Day 22, 4/16/09 -- Problems and solutions

Today started off with the gondola crew trying to figure out some oddness in the reaction wheel motor control. After lots of head-scratching, confusing tests, and emails across the pond, Michele finally noticed that the cables to the motor weren't plugged in all the way. The intermittent connection was enough to cause all sorts of non-repeatable weirdness.

This illustrates a point I often make: 90% of physics is trying to connect one thing to another. It could be a mechanical connection, electrical, thermal, data -- whatever. The hard part about experimental physics is always in the interfaces.

Jeff worked for most of today on his half-wave plate control system. I got started with the mounting hardware for the gondola protection stuff, though the gondola team was messing around with the gondola for much of the day so I didn't get a chance to drill some holes I need to drill. Perhaps tomorrow. Dan, who arrived last night from Minnesota, started setting up his calibration experiments.

The detector team is working hard on understanding the detectors we have operational. The fridges are still cold from yesterday, at about 260 mK. We had a bit of a snafu with our liquid helium order, though: When the truck was loaded, whoever loaded it neglected to secure our liquid helium dewars. Two of the three dewars fell over, the third was free to roll about on its casters in the back of the truck. Out of the 300L we were supposed to receive, we had only 28L delivered. Because we'll run out of helium in the cryostat on ~Sunday, probably, NASA is working hard to get us more by Saturday. Fingers crossed!

On the bright side, the NASA crew was working hard today on a very important improvement to the site: A new basketball hoop! Hopefully it'll be ready to use tomorrow.


No comments:

Post a Comment

Note: Only a member of this blog may post a comment.